# A Nonparametric Bayesian Local-Global Model for Enhanced Adverse Event Signal Detection in Spontaneous Reporting System Data

### v1

@@ 6 DISCUSSION
We propose nonparametric Bayesian models and their MCMC implementation for SRS data mining, balancing flexibility with shrinkage while capturing between-drug associations. Signal detection is performed through hypothesis tests that control FDR and minimize FNR, both estimated from MCMC posteriors. Our approach achieves significantly higher sensitivity than existing methods, particularly for rarer AEs and/or new drugs.
There are several limitations with our current approach that motivate interesting future directions. First, we focus here primarily on signal detection, i.e., to identify critical AEs of concern for each drug. The clusters of AEs produced by our DP priors have not been investigated but may reveal additional, deeper insights by identifying groups of “similar AEs”. A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups. Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI; Poleksic and Xie, 2019), possibly via appropriate prior distributions, may further enhance signal detection. Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis. Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
Acknowledgment. The authors are deeply grateful to Prof. Marianthi Markatou at the University at Buffalo for her insightful comments and helpful suggestions on the manuscript.
@@
This supplement provides a brief review of existing SRS data mining methods, additional details about the DP local-global models, detailed steps for the MCMC algorithms implementing our models, additional computation details, and additional simulation results. Sections, tables, and figures in this supplement are labeled with a prefix “S.”, such as, Section S.1, Table S.1, and Figure S.1, etc.
Appendix S.1 A REVIEW OF EXISTING BAYESIAN/EMPIRICAL BAYESIAN SRS DATA MINING METHODS AND THE ZIP MODEL-BASED LIKELIHOOD RATIO TEST METHOD
S.1.1 The GPS model
The GPS method (DuMouchel,, 1999) employs an empirical Bayes approach for stabilized signal detection within a Poisson model-based framework for the report counts $n_{ij}$. Specifically, we assume $n_{ij}\sim$ Poisson$(E_{ij}\lambda_{ij})$, with $\lambda_{ij}$ parametrizing the signal strength for the drug-$j$ and AE-$i$ pair. A two-component gamma mixture prior for $\lambda_{ij}$ is considered:
where $\kappa\in[0,1]$ is the mixing weight hyperparameter, and $\alpha_{1},\beta_{1}>0$, and $\alpha_{2},\beta_{2}>0$ are mixture component-specific gamma hyperparemeters. Here Gamma($\alpha$, $\beta$) denotes a probability distribution with density $f_{\text{gamma}}(y\mid\alpha,\beta)=\frac{1}{\Gamma(\alpha)}\beta^{\alpha}y^{\alpha-1}e^{-\beta y}$, $y>0$; $\alpha,\beta>0$. The posterior distribution of $\lambda_{ij}$ given $n_{ij}$, $E_{ij}$, and the hyperparameters is:
where $\kappa_{ij}=\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})/[\kappa f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{1},\beta_{1})+(1-\kappa)f_{\text{Gamma}}(\lambda_{ij}\mid\alpha_{2},\beta_{2})]$. Under an empirical Bayes approach, the point estimates $\hat{\kappa}$, $\hat{\alpha}_{1}$, $\hat{\beta}_{1}$, $\hat{\alpha}_{2}$, and $\hat{\beta}_{2}$ for the hyperparameters are obtained by maximizing their marginal likelihoods given $n_{ij}$ and $E_{ij}$. The resulting posterior of $\lambda_{ij}$ conditional on these estimated hyperparameters is used for inference and signal detection. Specifically, DuMouchel and Pregibon, (2001) define the empirical Bayesian score $E(\log_{2}(\lambda_{ij})\mid n_{ij},E_{ij},\hat{\kappa},\hat{\alpha}_{1},\hat{\beta}_{1},\hat{\alpha}_{2},\hat{\beta}_{2})$ as a point estimate of signal strength for the drug-$j$ and AE-$i$ pair, and also suggest determining the pair to be a signal if the 5th percentile (EB05) of the above conditional posterior for $\lambda_{ij}$ exceeds 2.
S.1.2 The BCPNN method
The BCPNN approach (Bate et al.,, 1998) assumes that conditional on the grand total count $n_{\bullet\bullet}$, the cell-specific counts $n_{ij}$, the marginal row totals $n_{i\bullet}$, and the marginal column totals $n_{\bullet j}$ each distributed as binomial-beta: (a) $n_{ij}\mid p_{ij}\sim\text{Binomial}(n_{\bullet\bullet},p_{ij}),\ p_{ij}\sim\text{Beta}(\alpha_{ij},\beta_{ij})$, (b) $n_{i\bullet}\mid p_{i\bullet}\sim\text{Binomial}(n_{\bullet\bullet},p_{i\bullet}),\allowbreak\ p_{i\bullet}\sim\text{Beta}(\alpha_{i\bullet},\beta_{i\bullet})$, (c) $n_{\bullet j}\mid p_{\bullet j}\sim\text{Binomial}(n_{\bullet\bullet},p_{\bullet j}),\allowbreak\ p_{\bullet j}\sim\text{Beta}(\alpha_{\bullet j},\beta_{\bullet j})$, where $p_{ij}$ is the cell reporting rate, and $p_{i\bullet}$ and $p_{\bullet j}$ are the marginal row and column reporting rates. (Note that the binomial model assumptions for $n_{ij}$, $n_{i\bullet}$ and $n_{\bullet j}$ can be understood as consequences of an independent Poisson model for $n_{ij}$ with certain choices of $p_{ij}$, $p_{i\bullet}$, and $p_{\bullet j}$.). Bate et al., (1998) consider the information content $\text{IC}_{ij}=\log_{2}\left[p_{ij}/(p_{i\bullet}p_{\bullet j})\right]$ as a measure of signal strength for the drug-$j$ AE-$i$ pair, with a higher IC indicating a stronger association. The authors suggest an empirical Bayes estimation of these quantities using $\hat{\alpha}_{ij}=\hat{\alpha}_{i\bullet}=\hat{\beta}_{i\bullet}=\hat{\alpha}_{\bullet j}=\hat{\beta}_{\bullet j}=1$ and $\hat{\beta}_{ij}=1/(E(p_{i\bullet}\mid n_{i\bullet})E(p_{\bullet j}\mid n_{\bullet j}))-1$, with the denominator of $\hat{\beta}_{ij}$ being computed using the posterior distributions $p_{i\bullet}\mid n_{i\bullet}\sim\text{Beta}(\hat{\alpha}_{i\bullet}+n_{i\bullet},\hat{\beta}_{i\bullet}+n_{\bullet\bullet}-n_{i\bullet})$ and $p_{\bullet j}\mid n_{\bullet j}\sim\text{Beta}(\hat{\alpha}_{\bullet j}+n_{\bullet j},\hat{\beta}_{\bullet j}+n_{\bullet\bullet}-n_{\bullet j})$. The consequent (empirical Bayes) posterior distribution of $\text{IC}_{ij}$ can be approximated using Monte Carlo simulation or an asymptotic normal distribution. A signal is detected at the drug-$j$ and AE-$i$ pair if the 2.5th posterior percentile for $\text{IC}_{ij}$ exceeds 0.
S.1.3 DP Poisson method of Hu et al., (2015)
Hu et al., (2015) consider the same likelihood $n_{ij}\sim$ Poisson($\lambda_{ij}E_{ij}$) as in GPS, but extend the two-gamma mixture prior for $\lambda_{ij}$ to a DP (infinite) mixture prior:
$\lambda_{ij}\mid(\alpha_{j},M_{0j})\sim\text{DP}(\alpha,M_{0j})=\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}$
where the base distribution $M_{0j}\equiv$ Gamma$(a_{j},a_{j})$ generates $\theta_{hj}\sim M_{0j}$ and the concentration hyperparameter $\alpha>0$ generates the mixture weights $\{\eta_{hj}\}_{h=1}^{\infty}$ satisfying $\sum_{h=1}^{\infty}\eta_{hj}=1$ with
$\eta_{1j}=v_{1j},\ \eta_{hj}=v_{hj}\prod_{t=1}^{h-1}(1-v_{tj}),\ h=2,\dots;\ v_{hj}\sim\text{independent Beta}(1,\alpha)\text{ for }h=1,\dots$ (5)
leveraging a stick-breaking representation (Sethuraman,, 1994) for each $j$. Hu et al., (2015) consider full Bayesian inference on $\lambda_{ij}$ by first eliciting a prior distribution on the hyperparameters $a_{j}$ and $\alpha_{j}$—specifically, independent vague priors $a_{j}\sim\text{Uniform}(0,1)$ and $\alpha_{j}\sim\text{Uniform}(0.2,10)$; $j=1,\dots J$—and then utilizing MCMC sampling from the resulting joint posterior. Using MCMC posterior draws for $\lambda_{ij}$, a signal at the drug-$j$ and AE-$i$ pair is identified if the 5th posterior percentile of $\lambda_{ij}$ exceeds a threshold of 2.
S.1.4 ZIP model based Pseudo LRT (frequentist) method
The likelihood ratio test based method and its variants (Huang et al.,, 2011, 2013, 2014) have an advantage in FDR control while maintaining high sensitivity. Additionally, the use of a post-hoc threshold makes it more suitable for practical applications. Consider the reporting count $n_{ij}\sim\text{ZIP}(n=n_{ij}\mid\lambda_{ij}\times E_{ij},\omega_{j})$, where $\lambda_{ij}$ is the relative reporting rate and $E_{ij}=n_{i\cdot}n_{\cdot j}/n_{\cdot\cdot}$ represents the expected count under independence. The pseudo LRT, as proposed by Chakraborty et al., (2022), is a recent method designed to accommodate the zero-inflated nature of SRS data while easing computational demands. Similar to the standard LRT method, this approach first tests a global alternative hypothesis $H_{aj}:\lambda_{ij}>1$ for at least one $i\in\{1,\dots,I\}$ against the null hypothesis $H_{0j}:\lambda_{ij}=1$ for all $i\in\{1,\dots,I\}$ for a given drug-$j$. If $H_{aj}$ is rejected, a step-down procedure is used to conduct individual tests on $H_{aij}:\lambda_{ij}>1$.
The ordinary LRT methods estimate the nuisance parameter $\omega_{j}$ from $H_{0j}$ and $H_{0j}\cup H_{aj}$ individually, the pseudo LRT is constructed using a global estimate $\hat{\omega}_{j}$ of $\omega_{j}$ consistent under $H_{0j}\cup H_{aj}$ in both the numerator and the denominator of the likelihood ratio test statistic. The resulting statistic is given by
$\text{LR}_{ij}=\exp(-(\hat{\lambda}_{ij}-1)E_{ij})\hat{\lambda}_{ij}^{n}.$
The maximum likelihood ratio statistic for drug-$j$ is then defined as $\text{MLR}_{j}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\}\end{subarray}}\text{LR}_{ij}$. In the global test, a parametric bootstrap can be used to approximate the p-value $P_{j}=\Pr(\text{MLR}_{j}\geq\text{MLR}_{j}^{\text{obs}})$ using (parametric bootstrap) draws for $\text{LR}_{ij}$ under $H_{0j}$ and hence $\text{MLR}_{j}$, where $\text{MLR}_{j}^{\text{obs}}$ is the observed test statistic. If the global test is rejected, the step-down p-values can be computed for the individual tests using the same parametric bootstrap samples for $\text{MLR}_{j}$ and using the individual observed likelihood ratio test statistic values, such that $P_{ij}=\Pr(\text{LR}_{ij}\geq\text{MLR}_{ij}^{\text{obs}})$. For multiple drugs, the same approach can be applied, defining an overall maximum likelihood ratio as $\text{MLR}=\max_{\begin{subarray}{c}i\in\{1,\dots,I\},j\in\{1,\dots,J\}\end{subarray}}\text{LR}_{ij}$. This step-down procedure avoids $\alpha$-inflation by using the same global test statistic, thus controlling the false discovery rate at the $\alpha$ level.
A consistent estimate of $\omega_{j}$ is obtained by maximizing the profile likelihood $\ell_{j}(\omega_{j})$, with $\hat{\lambda}_{ij}=\max\{1,n/E_{ij}\}$ as the MLE of $\lambda_{ij}$. The inference on zero-inflation is based on the estimate $\hat{\eta}_{ij}(\lambda_{ij},\hat{\omega}_{j})=\frac{\hat{\omega}_{j}}{\hat{\omega}_{j}+(1-\hat{\omega}_{j})\exp(-\lambda_{ij}E_{ij})}$, representing the estimated probability of zero-inflation for cells with zero count. A prediction of the zero-inflation indicator $z_{ij}$ can be made by generating Bernoulli samples from $\hat{\eta}_{ij}$.
@@
S.2.1 Complete hierarchical Bayesian representation of our models
@@
S.2.1.1 DP Poison local-global model
@@
[…]
@@
$v_{h}\sim\text{Beta}(1,\alpha)$
@@
S.2.1.2 DP ZIP local-global model
@@
[…]
@@
$D_{j}\equiv DP(\alpha_{j},G(\beta_{j}))\equiv\sum_{h=1}^{\infty}\eta_{hj}\delta_{\{\theta_{hj}\}}\approx\sum_{h=1}^{K}\eta_{hj}\delta_{\{\theta_{hj}\}}$
$\theta_{hj}\sim G_{0j}\equiv\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$
@@
$v_{hj}\sim\text{Beta}(1,\alpha_{j});v_{Kj}=1$
@@
$D\equiv DP(\alpha,G(\beta))\equiv\sum_{h=1}^{\infty}\eta_{h}\delta_{\{\theta_{h}\}}\approx\sum_{h=1}^{K}\eta_{h}\delta_{\{\theta_{h}\}}$
$\theta_{h}\sim G_{0}\equiv\text{Gamma}(\beta,\beta)$
@@
$v_{h}\sim\text{Beta}(1,\alpha)$
@@
S.2.2 Some notes on the “global-only” model
@@
In this section, we first provide MCMC algorithms implementing our models in Section S.3.1. The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes (Ishwaran and Zarepour,, 2002; Frühwirth-Schnatter and Malsiner-Walli,, 2019), and combine Gibbs sampling updates with stepping-out slice sampling updates (Neal,, 2003) for various parameters. Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.3.2. Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.3.3). The JAGS code (Plummer et al.,, 2003) for Hu et al., (2015)’s models is shown as a reference in Section S.1.3. The example statin dataset is included in Section S.3.5.
@@
    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution at $\theta$. $\pi(\beta)$ is the prior distribution of $\beta$. Direct random sampling from this density is not feasible. Instead, we use stepping-out slice sampling (Neal,, 2003) to generate Markov chain draws for $\beta$.
@@
    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
@@
S.3.2 Initial value generation for the MCMC algorithms
Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation, we initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$. The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution. The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set ${(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters. The probability weights $\eta_{hj}$ for these non-empty clusters are initialized using the relative frequencies of the $k$-means clusters, and the clusters are then relabeled to ensure $\eta_{1j}\geq\eta_{2j}\geq\dots\geq\eta_{K_{+}j}$. These atoms $\theta_{hj}$ are then used for numerical maximum likelihood estimation of $\beta$ with the model $\theta_{1j},\dots,\theta_{K_{+},j}\sim\text{independent~{}Gamma}(\text{shape}=\beta,\text{rate}=\beta)$, and the resulting estimate $\hat{\beta}$ is used to initialize $\beta_{j}$. The atoms $\theta_{h^{\prime}j}$ for the remaining empty clusters ($h^{\prime}=K_{+}+1,\dots,K$) are randomly drawn from $\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$ using the initialized $\beta_{j}$. Their corresponding weights $\eta_{h^{\prime}j}$ are initialized to $(\min_{h^{\prime}\in{1,\dots,K_{+}}}\eta_{h^{\prime}j})/10$, and all weights $\eta_{hj}$ for $h=1,\dots,K$ are subsequently renormalized to ensure $\sum_{h=1}^{K}\eta_{hj}=1$. The local signal strengths $\lambda_{ij}^{l}$ are then initialized for each $i$ using the cluster center (atom) corresponding to the $k$-means cluster containing point $i$ if $z_{ij}=1$, or by random sampling from ${\theta_{1j},\dots,\theta_{Kj}}$ with probabilities ${\eta_{1j},\dots,\eta_{Kj}}$. Finally, the concentration hyperparameter $\alpha_{j}$ is initialized by random sampling from a $\text{Gamma}(\text{shape}=1,\text{rate}=2)$ distribution.
@@
S.3.3 Optimization method in pFDR calculation
@@
For every quantile level in the $p$-grid, we obtain a matrix of test statistics $T_{ij}(p)$ based on the $p$-th posterior quantiles for $\lambda_{ij}$ computed from their MCMC draws. Subsequently, we compute FDR$(p,k)$ and FNR$(p,k)$ estimates using the formulas presented in the main text (Section 3.4) for all possible $p$ and $k$ defined by the grid. From the set of all $(p,k)$ pairs, we then retain only those pairs with computed FDR $\leq\alpha$ for a preset threshold $\alpha$ ($=0.05$ in our analyses). Finally, among the retained $(p,k)$, we identify the pair that minimizes FNR. If multiple $(p,k)$ yield the same minimum FNR, we select the pair with the smallest $p$ and the largest $k$.
S.3.4 JAGS code for DP Hu et al. model
@@
S.3.5 Example statin dataset
We present the reference SRS contingency table for statin drug in Table S.1. It is used for data generation in Section 4 and data analysis in Section 5.
Table S.1: The SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
@@
Acute Kidney Injury 1353 42 7 154 689 823 355651
Anuria 71 0 0 2 43 62 10403
Blood Calcium Decreased 14 2 0 0 110 17 15918
Blood Creatine Phosphokinase Abnormal 34 0 0 0 8 11 261
Blood Creatine Phosphokinase Increased 1175 125 32 200 562 768 23805
Blood Creatine Phosphokinase Mm Increased 2 0 0 0 9 0 14
Blood Creatinine Abnormal 27 0 0 0 5 3 3385
Blood Creatinine Increased 227 10 0 17 210 97 74742
Chromaturia 340 10 6 33 174 114 19294
Chronic Kidney Disease 152 16 2 19 177 37 339179
Compartment Syndrome 53 0 0 1 12 12 2644
Creatinine Renal Clearance Decreased 6 0 0 2 124 6 7768
Diaphragm Muscle Weakness 14 0 0 0 8 1 94
Electromyogram Abnormal 2 0 0 2 0 0 132
End Stage Renal Disease 30 0 0 0 19 6 97553
Glomerular Filtration Rate Abnormal 8 0 0 0 1 0 1069
Glomerular Filtration Rate Decreased 59 1 0 8 39 29 13190
Hypercreatininaemia 0 0 0 0 8 0 648
Hypocalcaemia 36 0 0 16 8 18 23102
Muscle Disorder 291 2 7 21 191 87 7329
Muscle Enzyme Increased 48 1 0 0 13 9 410
Muscle Fatigue 85 0 2 16 30 39 4257
Muscle Haemorrhage 24 0 0 5 13 4 3806
Muscle Necrosis 68 2 0 1 10 20 662
Muscle Rupture 181 25 0 61 36 120 3219
Muscular Weakness 1857 45 31 152 808 859 111003
Musculoskeletal Discomfort 137 18 15 25 187 93 19931
Musculoskeletal Disorder 56 3 0 9 65 73 25881
Musculoskeletal Pain 420 3 2 38 324 228 82576
Myalgia 5362 341 151 939 2757 3216 143819
Myasthenic Syndrome 1 0 9 0 3 6 643
Myoglobin Blood Increased 71 4 0 4 16 39 1003
Myoglobin Blood Present 2 0 0 0 0 0 8
Myoglobin Urine Present 0 0 0 0 1 2 70
Myoglobinaemia 15 0 0 0 0 0 62
Myoglobinuria 26 4 0 1 1 9 296
Myopathy 849 64 45 145 377 544 6695
Myopathy Toxic 31 0 0 1 2 21 457
Myositis 219 8 10 28 62 141 7482
Necrotising Myositis 279 0 0 2 10 52 278
Oliguria 52 0 0 4 24 37 7590
Renal Failure 534 26 11 69 225 195 250710
Renal Impairment 390 52 11 37 161 169 103343
Renal Tubular Necrosis 40 0 0 12 10 24 12762
Rhabdomyolysis 2041 52 44 163 936 1376 31707
Tendon Discomfort 9 0 0 3 22 10 794
Other Pt 180699 4886 2845 20296 113960 76068 61724222
@@
S.4.1 Simulation results for DP-based models

### v2

@@ 7 DISCUSSION
We propose nonparametric Bayesian models and their MCMC implementation for SRS data mining, balancing flexibility with shrinkage while capturing between-drug associations. Signal detection is performed through hypothesis tests that control FDR and minimize FNR, both estimated from MCMC posteriors. Our approach achieves significantly higher sensitivity than existing methods, particularly for rare AEs and/or newer drugs, with small observed counts. Application to the statin dataset demonstrates our model’s superior performance, which detects more plausible signals involving rare AEs and/or newer drugs than competing methods while reliably identifying well-established AEs. By enabling between-drug information sharing, our model can also discover potential signals among rare AEs and provide more nuanced detection for semi-rare AEs.
There are several limitations with our current approach that motivate interesting future directions. First, we focus here primarily on signal detection, i.e., to identify critical AEs of concern for each drug. The clusters of AEs produced by our DP priors have not been investigated but may reveal additional, deeper insights by identifying groups of “similar AEs”. A reformulation via mixed membership models could further enhance interpretability by allowing the same AE to appear in multiple clusters/groups. Second, while we consider data-driven characterization of similarities in AE signal strengths between drugs, a more flexible approach acknowledging known Drug-Drug Interactions (DDI poleksic2019database ), possibly via appropriate prior distributions, may further enhance signal detection. Finally, our current framework aggregates drug-AE data across time points for a cross-sectional analysis. Extending the model to accommodate temporal variations in SRS data could enable formal identification of evolving AE patterns.
Acknowledgment. The authors are deeply grateful to Prof. Marianthi Markatou at the University at Buffalo for her insightful comments and helpful suggestions on an earlier version of the manuscript.
@@
This supplement provides additional details about the DP local-global models, detailed steps for the MCMC algorithms implementing our models, additional computation details, and additional simulation results. Sections, tables, and figures in this supplement are labeled with a prefix “S.”, such as, Section S.1, Table S.1, and Figure S.1, etc.
@@
S.1.1 Complete hierarchical Bayesian representation of our models
@@
S.1.1.1 DP Poison local-global model
@@
$\text{DP local-global prior for a non-baseline drug }j\in\{1,\dots,J-1\}\!:$
@@
$v_{h}\sim\text{Beta}(1,\alpha),h=1,\dots,K-1;\ v_{K}=1$
@@
S.1.1.2 DP ZIP local-global model
@@
$\text{DP local-global prior for a non-baseline drug }j\in\{1,\dots,J-1\}\!:$
@@
$v_{hj}\sim\text{Beta}(1,\alpha_{j}),h=1,\dots,K-1;\ v_{Kj}=1$
@@
$v_{h}\sim\text{Beta}(1,\alpha),h=1,\dots,K-1;\ v_{K}=1$
@@
S.1.2 Some notes on the “global-only” model
@@
In this section, we first provide MCMC algorithms implementing our models in Section S.2.1. The MCMC algorithms are derived using finite mixture approximation of the underlying DP processes [23, 16], and combine Gibbs sampling updates with stepping-out slice sampling updates [31] for various parameters. Some additional notes on the MCMC implementation, including initial value generation for our MCMC algorithms are then provided in Section S.2.2. Next we provide detailed notes on the grid-search method identifying optimal quantile-level $p$ and threshold $k$ used for our hypothesis tests for signal detection (Section S.2.3). The JAGS code [32] for Hu et al.’s model [18] is shown as a reference in Section S.2.4. The example statin dataset is included in Section S.2.5.
@@
    Here $f_{\text{Gamma}}(\theta\mid\text{shape}=\beta,\text{rate}=\beta)$ denotes the probability density function of the Gamma($\beta,\beta$) distribution at $\theta$. $\pi(\beta)$ is the prior distribution of $\beta$. Direct random sampling from this density is not feasible. Instead, we use stepping-out slice sampling [31] to generate Markov chain draws for $\beta$.
@@
    $p_{ij}^{l}=\frac{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})}{\pi^{l}f_{\text{Poisson}}(n_{ij}\mid\lambda_{ij}^{l}E_{ij})+(1-\pi^{l})f_{\text{Poisson}}(n_{ij}\mid\lambda_{i}^{g}E_{ij})}.$
@@
S.2.2 Initial value generation for the MCMC algorithms
Given the observed data $n_{ij}$ and $E_{ij}$ for $i=1,\dots,I$, $j=1,\dots,J$, and a prespecified number $K$ of DP mixture components for finite approximation. The default value for $K$ in our software is set to $\max(\log(IJ),10)$, a setting that is generally sufficient for most cases. For the simulation study in Section 5, we used $K=10$, which satisfies the condition $K>\log(1492\times 7)$. In the data analysis presented in Section 6, we chose a more conservative value of $K=20$ to allow for a larger number of atoms in the DP, facilitating a more cautious interpretation of the results. We initialize the number of nonempty components $K_{+}$ to the smallest integer greater than $\min\{K/2,I-1\}$. The local-global indicators $z_{ij}^{l}$ are randomly initialized by independently sampling each cell $(i,j)$ from a $\text{Bernoulli}(0.5)$ distribution. The DP parameters for each local and global DP process are then initialized using the subset of data $(n_{ij},E_{ij})$ assigned to that process, as determined by the indicators $z_{ij}^{l}$, for $i=1,\dots,J-1$.
For the initialization of parameters in the local DP prior for the $j$-th drug ($j=1,\dots,J-1$), we perform $k$-means clustering with the initialized $K_{+}$ number of clusters on the set $\{(n_{ij}+0.5)/(E_{ij}+0.5):z_{ij}=1,\ i=1,\dots,I\}$, and use the resulting cluster centers as initial values for the atoms $\theta_{hj}$, $h=1,\dots,K_{+}$, of the non-empty clusters. The probability weights $\eta_{hj}$ for these non-empty clusters are initialized using the relative frequencies of the $k$-means clusters, and the clusters are then relabeled to ensure $\eta_{1j}\geq\eta_{2j}\geq\dots\geq\eta_{K_{+}j}$. These atoms $\theta_{hj}$ are then used for numerical maximum likelihood estimation of $\beta$ with the model $\theta_{1j},\dots,\theta_{K_{+},j}\sim\text{independent~Gamma}(\text{shape}=\beta,\text{rate}=\beta)$, and the resulting estimate $\hat{\beta}$ is used to initialize $\beta_{j}$. The atoms $\theta_{h^{\prime}j}$ for the remaining empty clusters ($h^{\prime}=K_{+}+1,\dots,K$) are randomly drawn from $\text{Gamma}(\text{shape}=\beta_{j},\text{rate}=\beta_{j})$ using the initialized $\beta_{j}$. Their corresponding weights $\eta_{h^{\prime}j}$ are initialized to $(\min_{h^{\prime}\in{1,\dots,K_{+}}}\eta_{h^{\prime}j})/10$, and all weights $\eta_{hj}$ for $h=1,\dots,K$ are subsequently renormalized to ensure $\sum_{h=1}^{K}\eta_{hj}=1$. The local signal strengths $\lambda_{ij}^{l}$ are then initialized for each $i$ using the cluster center (atom) corresponding to the $k$-means cluster containing point $i$ if $z_{ij}=1$, or by random sampling from $\theta_{1j},\dots,\theta_{Kj}$} with probabilities $\eta_{1j},\dots,\eta_{Kj}$}. Finally, the concentration hyperparameter $\alpha_{j}$ is initialized by random sampling from a $\text{Gamma}(\text{shape}=1,\text{rate}=2)$ distribution.
@@
S.2.3 Optimization method in pFDR calculation
@@
For every quantile level in the $p$-grid, we obtain a matrix of test statistics $T_{ij}(p)$ based on the $p$-th posterior quantiles for $\lambda_{ij}$ computed from their MCMC draws. Subsequently, we compute FDR(p,k)(p,k) and FNR(p,k)(p,k) estimates using the formulas presented in the main text (Section 4.4) for all possible $p$ and $k$ defined by the grid. From the set of all $(p,k)$ pairs, we then retain only those pairs with computed FDR $\leq\alpha$ for a preset threshold $\alpha$ (=0.05=0.05 in our analyses). Finally, among the retained $(p,k)$, we identify the pair that minimizes FNR. If multiple $(p,k)$ yield the same minimum FNR, we select the pair with the smallest $p$ and the largest $k$.
S.2.4 JAGS code for DP Hu et al. model
@@
S.2.5 Example statin dataset
We present the first 30 rows of the reference SRS contingency table for statin drug in Table S.1. It is used for data generation in Section 5 and data analysis in Section 6.
Table S.1: The first 30 rows of the SRS contingency table summarized from the FDA FEARS database (Q3 2014-Q4 2020) which catalogs a total of 63,976,610 reported cases.
@@
Abasia 117 0 7 9 84 45 18046
Abdominal Adhesions 10 0 0 0 3 2 3831
Abdominal Discomfort 501 33 14 92 497 217 214807
Abdominal Distension 267 7 5 57 251 262 101261
Abdominal Pain 655 27 7 116 428 372 232708
Abdominal Pain Lower 60 1 1 2 52 22 22527
Abdominal Pain Upper 696 14 15 78 554 266 197925
Abdominal Rigidity 12 0 0 1 1 12 2919
Abdominal Symptom 7 0 0 0 2 2 718
Abdominal Tenderness 20 0 0 0 5 10 5783
Abnormal Behaviour 72 2 5 9 23 77 35802
Abnormal Faeces 23 0 0 1 18 8 7982
Abnormal Loss Of Weight 89 0 1 3 15 79 9042
Abnormal Sensation In Eye 4 0 0 0 3 0 2571
Abortion Induced 10 0 0 3 1 55 6668
Abortion Spontaneous 32 2 0 5 12 4 28245
Abscess Limb 4 0 0 0 4 0 5258
Accident 24 0 0 9 28 6 12034
Accident At Work 0 0 0 0 14 2 2087
Accidental Exposure To Product 13 2 1 3 12 7 46642
Accidental Overdose 56 0 0 11 7 36 42543
Accidental Underdose 4 0 0 0 2 0 7423
Acidosis 55 3 0 4 4 22 8117
Acne 28 0 1 10 35 22 105646
Activities Of Daily Living Impaired 135 1 3 15 82 200 14483
Acute Coronary Syndrome 105 1 0 1 66 63 9259
Acute Generalised Exanthematous Pustulosis 30 0 0 0 11 3 8161
Acute Hepatic Failure 107 0 10 1 18 66 14260
@@
S.3.1 Simulation results for DP-based models